# Addressing Source Scale Bias via Image Warping for Domain Adaptation

### v1

@@ Abstract
In visual recognition, scale bias is a key challenge due to the imbalance of object and image size distribution inherent in real scene datasets. Conventional solutions involve injecting scale invariance priors, oversampling the dataset at different scales during training, or adjusting scale at inference. While these strategies mitigate scale bias to some extent, their ability to adapt across diverse datasets is limited. Besides, they increase computational load during training and latency during inference. In this work, we use adaptive attentional processing – oversampling salient object regions by warping images in-place during training. Discovering that shifting the source scale distribution improves backbone features, we developed a instance-level warping guidance aimed at object region sampling to mitigate source scale bias in domain adaptation. Our approach improves adaptation across geographies, lighting and weather conditions, is agnostic to the task, domain adaptation algorithm, saliency guidance, and underlying model architecture. Highlights include +6.1 mAP50 for BDD100K Clear DENSE Foggy, +3.7 mAP50 for BDD100K Day Night, +3.0 mAP50 for BDD100K Clear Rainy, and +6.3 mIoU for Cityscapes ACDC. Our approach adds minimal memory during training and has no additional latency at inference time. Please see Appendix for more results and analysis.
![[Uncaptioned image]](x1.png)
@@ 1 Introduction
Visual detection and segmentation of things (i.e. object instances) and stuff, across different scales is a fundamental challenge in computer vision [49, 35, 19, 45]. Natural scene datasets used for training different tasks often exhibit bias of imaged objects towards certain scales (or sizes). For instance, in autonomous driving situations, camera perspective and field of view warrant that many more distant objects that appear small in the image are captured compared to nearby objects that appear large, even if the real object 3D scale distribution is near-uniform. This results in an inherent imbalance or bias in the distribution of sizes of object projections in the image, leading to sub-optimal recognition models.
Common approaches to address scale bias are: (a) selecting an image subset to create a uniform scale distribution but this is suboptimal as only a fraction of the available data is used; (b) oversampling via data augmentation (e.g., multi-scale training [35] and multi-scale inference [10, 34, 13]), while helpful, also introduces significant computational and storage overhead; (c) introducing scale-invariance priors (e.g., feature pyramids [21, 23, 22]), but this only partially alleviates scale bias, as both things and stuff vary widely in scale and image locations.
We propose a new method to address scale bias, orthogonal to previous techniques, using adaptive attentional processing. We warp images in-place while training models to shift the unbalanced scale distribution of objects (see Figure 2). While non-uniform warping techniques, i.e. allocating more pixels using a predefined guidance, have been applied to improve image classification [25], detection [36, 9, 37, 7] and segmentation [37] in supervised settings, they do not principally examine and address the scale bias for domain adaptation.
We propose Instance-level Saliency Guidance which analytically shifts the object scale distribution to remove bias or match target distribution by leveraging ground truth labels for maximum performance gain. Because the warping is in-place, we directly re-sample objects (e.g. expand smaller ones) and squeeze the large empty regions in the scene (e.g. road), allowing for a better use of the image pixel budget for balanced distribution training. We show that this warping improves backbone features of our model crucial for domain adaptation. To make sure our outputs are in the original unwarped configuration, we unwarp the features, as in LZU [37], and train without warping the object labels. During inference, no warping is performed and the method requires no additional computational latency. Our contributions are highlighted as follows:
We are the first work to show that scale bias in real-to-real domain adaptation can be effectively addressed through image warping, and that shifting source scale distribution significantly improves backbone features.
@@
Our approach is agnostic to task (detection or segmentation), adaptation strategy, model architecture and warping guidance used for sampling.
Our approach improves domain adaptation across various lighting (+3.7 mAP50 for BDD100K Day Night detection), weather conditions (+6.1 mAP50 for BDD100K Clear DENSE Foggy detection), and geographies (+6.3 mIoU for Cityscapes ACDC semantic segmentation). See Figure 1.
@@
Our approach is efficient: we introduce train-time image warping and feature unwarping for domain adaptation without need for test-time warping. This results in minimal training overhead and no additional inference latency.
@@ 2 Related Work
Alleviating Scale Bias. Recognition across different scales is a historically important problem. Priors that encourage scale invariance like feature (or anchor) pyramids are widely used in pre-deep-learning approaches [23, 5] and recently [27, 21, 22]. While helpful, these priors do not solve scale bias completely, as an object can be bigger or smaller than pre-defined scales. Common augmentation techniques such as multi-scale training [27, 35, 17, 13] can considerably increase training time. Similarly, test time scaling [10, 34] and multi-scale or multi-crop predictions [26, 13] improve accuracy at the cost of latency and memory [15]. Another line of works select “optimal” scale at inference [3, 8], but don’t improve the learned model itself. Compared to earlier work, our method explicitly addresses scale bias requiring (a) little additional memory during training since we warp images in-place; and (b) no additional inference latency as we do not warp during testing. Crucially, our approach is orthogonal to other scale bias reduction techniques, and can be used in a plug-and-play manner.
Guided Image Warping. Non-uniform Spatial Transformations [42] have been used to correct image distortions arising from camera acquisition. Recently, spatial transformation methods [25, 16] for learned models have advocated introducing image distortions instead. Learning to Zoom [25] showed over-sampling task-specific salient regions is useful as visual recognition requires combining information from different spatial resolutions. Building on this paradigm, warping for other tasks like detection [36, 37] and segmentation [37] were explored. Fovea [36], and Two Plane Prior [9] designed “salient” priors to guide warping, introducing static & temporal, and geometric saliency guidance respectively.
Our approach stands outs in the following ways: (a) Unlike previous image warping efforts focusing on supervised settings, our innovative training strategy and saliency guidance tailor image warping for unsupervised domain adaptation to directly address scale bias. (b) Unlike former image warping methods that indiscriminately process objects and background elements, we introduces an analytical instance-level saliency guidance that targets regions containing objects.
Image Transformations for Domain Adaptation. Image Transformations like flipping, cropping, blurring, color jittering, grayscaling, contrast adjustment are widely used as perturbations for domain adaptation [39, 45, 46, 24, 31, 48, 38, 6, 41, 43, 18, 12, 13, 11, 14, 17] to increase source domain diversity and decrease domain gap to the target domain. For example, 2PCNet [17] introduces a domain-specific Night Augmentation (NightAug) pipeline that attempts to make day images appear closer to night images. To the best of our knowledge, in-place image warping as a domain adaptive perturbation has not been explored in prior works. Our approach is agnostic to the specific domain adaptation training algorithm, and can be easily incorporated with any other augmentation strategy.
@@ 3 Method
We first describe how we warp images assuming a spatial guidance function while remaining task agnostic. Then, we design our warping guidance that can explicitly address scale bias (as shown in Figure 1). Finally, we detail how to train models with warped images that adapt better to new domains. The entire strategy is illustrated in Figure 3.
@@ 3.1 Image Warping and Feature Unwarping
We employ the saliency guided sampling mechanism from [25]. However, other methods may also be used (like thin plate spline transformations [16, 7]).
Image Warping: We consider an inverse transformation parameterized by saliency such that,
@@
Feature Unwarping: If we operate in the space of warped images, our outputs would also be warped. However, we do not wish to warp our labels to match the input warped image, hence, we perform feature unwarping of spatial backbone features and train the model without warping our labels or unwarping our predictions, like [36, 9]. The benefit of this strategy is that it renders our warping mechanism task agnostic. We adopt the strategy from LZU [37], which approximates with , a piecewise tiling of invertible bilinear maps, and then perform feature unwarping via,
We restrict the space of allowable warps for such that axis alignment is preserved. Entire rows or columns are “stretched” or “compressed” in this case, which has been observed to perform better [36, 9, 37]:
where is a Gaussian kernel and denote the two image axes in the pixel space. and are marginalized versions of along and respectively.
@@ 3.2 Designing Saliency Guidance
Warping guidance decides which image regions are warped (e.g., scaled) more than others. Since images are warped in-place, if some regions are expanded, others must be reduced in size (Figure 4). For the warping mechanism considered [25, 36, 37], this requires designing a saliency function . We investigate three saliency functions, based on two approximate guidances proposed in previous work [36, 37, 9] and one that we propose in this work.
Approximate Saliency Guidance: The Geometric Prior [9] assumes that a dominant vanishing point exists in the distant scene, and designs a saliency map to exploit this property and oversample objects that appear small. But this prior tries to warp the entire scene that is farther away and the objects themselves may not be oversampled sufficiently to overcome size bias (Figure 4).
The Static Prior [36, 37] assumes that training set statistics will remain valid at test time, which is often not the case in domain adaptation due to domain shift and label unavailability. This incorrect assumption leads to the creation of a saliency map targeting average object locations, which may not accurately represent specific target images where objects deviate from these averages (Figure 4). Moreover, warping during the test phase, after adaptation, increases inference latency and is less effective due to the unknown scale distribution of the target domain (See Section 5.4 for ablation). Essentially, these saliency approaches do not utilize actual object labels and fail to concentrate on oversampling the objects, a strategy that would significantly benefit the shifting of the object scale distribution.
Instance-level Saliency Guidance. We propose Instance-level Saliency Guidance which explicitly shifts the object scale distribution by leveraging ground truth labels (Figure 4). During training, source image has corresponding labels available. For simplicity, we assume the labels are bounding boxes.
To shift the object scale distribution, we oversample image regions corresponding to every object instance. Following [1, 36], we use kernel density estimation to model the bounding box saliency as a sum of Gaussians,
@@
We propose an expansion factor to decide the warping extent of all objects,
@@
where is the fraction of object size corresponding to an interval between size thresholds and (See Figure 1). In our case, we consider , to represent the percentage of the small, medium, and large objects in the dataset, defined by COCO [20] thresholds for object sizes in the image.
@@
Saliency scale and expansion factor are inversely proportional, thus where is a constant. We clip between of 0 and , since negative saliency values lead to holes, and an upper bound to discourage extreme warping.
When segmentation masks are available but not bounding boxes, we convert them to boxes, termed “from-seg” boxes. This conversion minimally impacts performance, as detailed in our ablations in Section 5.4 and Supplementary.
3.3 Warping for Domain Adaptation
@@
We consider unsupervised domain adaptation, wherein source domain is available with labels, i.e. dataset and target domain only consists of images, .
@@
Domain Adaption Framework: The standard framework comprises of two stages. First, a supervised pre-training stage where a model is trained using supervised loss on source domain, Second, in the unsupervised adaptation stage the model is adapted to target domain using dataset through an additional unsupervised loss. Generally, a student and teacher (initialized using ) are trained jointly through self-training, by employing psuedo labels [17, 12]. While training the student, a supervised loss is used to preserve performance on the source domain in a supervised branch:
@@
The teacher is updated as the exponential moving average (EMA) of the student.
Warping and Unwarping in the Domain Adaptation Framework: The design decisions that guide our choices are: (a) Maximize transferability of features from supervised stage to unsupervised training stages. (b) Warping at test time after adaptation is likely to be suboptimal as the target domain’s scale distribution is unknown and can introduce latency. If the target distribution is known, the Instance-level guidance can be shifted analytically to match this distribution. Thus, our warping mechanism operates only where images from source domain are used, i.e. it is employed in the supervised pre-training stage and the student model in the unsupervised stage when source image is forwarded. We do experiment by warping target images too while employing Static [36, 37] and Geometric prior [9] saliency guidances (which do not use ground truth object labels to generate saliency), however, that reduces performance (See Table 8).
Consider the recognition model as a combination of backbone and prediction heads , i.e. . We warp the image and then unwarp the feature maps from backbone respectively to compose,
@@
following [37] for a source image . Thus, losses are modified for the supervised pre-training stage as,
and for the unsupervised training stage as,
@@
For our experiments, we use DAFormer [12] and 2PCNet [17] as our baseline segmentation and detection domain adaption algorithms respectively. We do not make additional assumptions about the domain adaptation setup, and thus our method is agnostic to the task specification (segmentation or detection). Other domain adaption algorithms [13, 14, 18, 6, 48] are also applicable.
@@
Inference: We employ the teacher model during inference, and do not require any image warping and feature unwarping at test time. This is in contrast with image warping techniques proposed in the supervised setting [36, 9, 37], which incur overheads from test-time warping and assume additional information (like vanishing points [9]) to generate saliency maps at test time.
@@
Supervised Setting: Our training method can also be applied in supervised learning scenario by following the supervised pre-training steps to train a model . That is, omitting the unsupervised training stage completely and only performing supervised pre-training stage – then removing the warping and unwarping layers completely from the trained model. This is unlike prior work [25, 36, 8, 37] which advocates warping at test time too. Model performance improves significantly on target domain without performing adaptation (See Table 5). For Source Domain Results, see Supplementary Section C.
4 Experimental Methodology and Datasets
We implement our source domain warping technique alongside a baseline domain adaptation algorithm, ensuring fair comparisons by using the same model, datasets, training schedules, hyperparameters, and random seeds.
Object Detection: Following the recent and popular domain adaptation strategy 2PCNet [17] (2023), we use Faster R-CNN [27] with ResNet-50 [10] adhering to their training hyperparameters and protocols. While 2PCNet [17] solely focuses on Day-to-Night adaptation, our method spans various adaptation scenarios. For scenarios other than day-to-night adaptation, we exclude the NightAug augmentation proposed in [17] for both the baseline and our method.
Semantic Segmentation: Aligning with the recent and popular domain adaptation strategy DAFormer [12] (2022), we employ the same SegFormer [44] head and MiT-B5 [44] backbone, and follow their training hyperparameters, protocols, and seed for fair comparison. While DAFormer [12] shows results on Sim2Real Gap which is not a focus of our work, results and discussion are available in Supplementary Section A. Extensions of DAFormer, including HRDA [13] and MIC [14], also benefits from our methodology, detailed in the Appendix B
4.1 Datasets
BDD100K [47] features 100,000 images with a resolution of 1280x720 for object detection and segmentation, covering various weather conditions and times of day. It includes 10 categories for annotation.
Cityscapes [4] provides 5,000 images of urban road scenes at a resolution of 2048x1024 with clear weather for semantic segmentation, with 19 categories.
ACDC [32] is designed for adverse conditions (fog and snow), including 1,600 images at 2048x1024 for segmentation across 19 categories.
DarkZurich [30] is tailored for low-light conditions, offering 2,416 unlabeled nighttime images and 151 labeled twilight images for segmentation, with a resolution of 1920x1080, and a focus on urban settings.
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl | person | car | truck | bike |
2PCNet | 27.0 | 45.0 | 30.1 | 0.0 | 15.4 | 35.5 | 87.5 | 72.5 | 19.1 | 0.9 |
+ Ours w/ Sta. Prior | 28.6 | 48.5 | 31.7 | 0.0 | 19.6 | 34.0 | 88.1 | 72.5 | 25.6 | 7.7 |
+ Ours w/ Geo. Prior | 27.1 | 45.8 | 29.5 | 0.0 | 15.2 | 37.5 | 86.5 | 75.4 | 21.0 | 0.4 |
+ Ours w/ Inst. Warp | 33.1 | 51.1 | 39.7 | 0.0 | 14.5 | 45.2 | 85.8 | 66.3 | 50.0 | 2.2 |
bike |
2PCNet | 25.8 | 48.4 | 24.4 | 10.9 | 34.1 | 49.0 | 56.3 | 40.4 | 77.0 | 57.2 | 55.9 | 36.8 | 36.8 | 55.6 | 68.1 | ||||||
+ Ours w/ Sta. Prior | 26.8 | 50.6 | 24.9 | 11.4 | 35.9 | 49.6 | 57.4 | 42.5 | 77.4 | 60.1 | 59.5 | 44.8 | 39.4 | 55.7 | 68.9 | ||||||
+ Ours w/ Geo. Prior | 26.9 | 49.9 | 24.5 | 11.8 | 35.6 | 50.1 | 56.1 | 44.6 | 77.5 | 58.0 | 58.6 | 33.3 | 45.8 | 55.8 | 69.5 | ||||||
+ Ours w/ Inst. Warp | 27.1 | 51.4 | 24.6 | 11.0 | 36.7 | 49.0 | 57.2 | 46.6 | 77.5 | 59.8 | 60.9 | 46.3 | 42.2 | 54.9 | 68.3 |
bike |
2PCNet | 23.5 | 46.4 | 21.1 | 9.5 | 26.0 | 43.3 | 54.4 | 30.8 | 73.1 | 53.8 | 55.2 | 37.5 | 44.5 | 49.4 | 65.2 | ||||||
+ Ours w/ Sta. Prior | 25.3 | 49.4 | 22.4 | 10.3 | 27.4 | 45.1 | 58.2 | 39.3 | 75.0 | 53.6 | 54.5 | 41.4 | 49.6 | 52.9 | 69.6 | ||||||
+ Ours w/ Geo. Prior | 25.6 | 49.7 | 22.9 | 10.0 | 27.7 | 45.7 | 58.8 | 41.8 | 75.0 | 55.8 | 55.2 | 40.8 | 49.3 | 50.5 | 69.4 | ||||||
+ Ours w/ Inst. Warp | 25.6 | 50.1 | 22.6 | 9.8 | 27.6 | 46.6 | 58.9 | 42.7 | 75.1 | 55.0 | 55.5 | 42.6 | 50.1 | 51.9 | 69.0 |
bike | ||
2PCNet | 16.6 | 30.8 | 14.9 | 6.0 | 24.8 | 26.6 | 41.0 | 20.0 | 77.0 | 41.8 | 17.1 | 29.9 | 19.3 | ||
+ Ours w/ Sta. Prior | 17.3 | 31.0 | 16.7 | 6.2 | 25.9 | 29.5 | 42.2 | 20.9 | 76.7 | 40.5 | 21.7 | 24.4 | 21.8 | ||
+ Ours w/ Geo. Prior | 16.6 | 30.3 | 15.4 | 5.2 | 25.1 | 28.7 | 42.3 | 19.2 | 75.0 | 39.5 | 18.1 | 26.7 | 21.3 | ||
+ Ours w/ Inst. Warp | 17.9 | 32.1 | 18.2 | 6.8 | 26.7 | 32.1 | 45.7 | 20.6 | 77.6 | 43.2 | 20.8 | 26.7 | 22.2 |
5 Experiments

### v2

# Saliency Guided Image Warping for Unsupervised Domain Adaptation

@@ Abstract
Driving is challenging in conditions like night, rain, and snow. The lack of good labeled datasets has hampered progress in scene understanding under such conditions. Unsupervised domain adaptation (UDA) using large labeled clear-day datasets is a promising research direction in such cases. Current UDA methods, however, treat all image pixels uniformly, leading to over-reliance on the dominant scene backgrounds (e.g., roads, sky, sidewalks) that appear dramatically different across domains. As a result, they struggle to learn effective features of smaller and often sparse foreground objects (e.g., people, vehicles, signs).
In this work, we improve UDA training by using in-place image warping to focus on salient object regions. Our insight is that while backgrounds vary significantly across domains (e.g., snowy night vs. clear day), object appearances vary to a lesser extent. Therefore, we design instance-level saliency guidance to adaptively oversample object regions, which reduces adverse effects from background context and enhances backbone feature learning. We then unwarp the better learned features while adapting from source to target. Our approach improves adaptation across geographies, lighting, and weather conditions, and is agnostic to the task (segmentation, detection), domain adaptation algorithm, saliency guidance, and underlying model architecture. Result highlights include +6.1 mAP50 for BDD100K Clear DENSE Foggy, +3.7 mAP50 for BDD100K Day Night, +3.0 mAP50 for BDD100K Clear Rainy, and +6.3 mIoU for Cityscapes ACDC. Our method adds minimal training memory and incurs no additional inference latency. Please see Appendix for more results and analysis.
@@ 1 Introduction
Consider driving in safety-critical scenarios such as at night, in rain, or snow. Robust detection and segmentation of objects (e.g., cars and pedestrians) and background are critical to prevent accidents. In these conditions, human perception focuses on salient objects rather than background for decision making [39], using high-level semantics to search and decide where to focus [10].
Camera-captured images, meanwhile, exhibit an object-background pixel imbalance, where pixel distribution is skewed towards backgrounds taking up more space. In contrast, objects are smaller (e.g. traffic signs), more scattered, and thus occupy fewer pixels. The camera’s perspective further amplifies this effect, making distant objects appear smaller and allowing larger backgrounds to dominate the scene. However, contemporary perception models treat image pixels, whether objects or backgrounds, uniformly. This leads to an over-reliance on the dominant backgrounds, making it challenging to learn effective features for less prevalent objects amid these dominant backgrounds.
This over-reliance on background pixels becomes even more problematic in Unsupervised Domain Adaptation (UDA) [13, 14, 17], which is crucial for driving in challenging conditions due to the scarcity of labeled data. However, UDA exhibits differences in cross-domain object-background variations, where background regions show greater variations than object regions across domains. Consequently, this over-reliance on background pixels makes adaptation difficult. Reducing dependence on these variable backgrounds would improve adaptation to new domains.
One possible remedy is to train at various image scales using data augmentation (e.g., multi-scale training [33] and multi-scale inference [11, 32, 14]). However, these strategies introduce significant computational and storage overhead. Another approach, employing scale invariance priors (e.g., feature pyramids [19, 22, 21] or test-time scale optimization [8, 4]), only partially improve recognition of different sizes of objects and introduces additional inference latency. Yet another strategy is to oversample or ‘zoom’ into salient image regions [24, 35, 7, 9, 36]. However, these methods do not distinguish between object and background, are only applicable to supervised settings, and introduce additional inference latency due to test-time operations.
To address these limitations, we develop a novel strategy that judiciously utilizes the available pixel space in an image to focus on salient regions for unsupervised domain adaptation. We adaptively oversample salient image regions corresponding to objects in-place, directly warping the original image without creating new images or additional crops. Then, we unwarp the better learned features while adapting models from a source to a target domain. This reduces background context dependence for detection and segmentation, significantly improving domain adaptability. During inference, we do not perform any warping as the learned features are already improved, thus adding no latency. We show significant improvements on a large set of real to real adaptation scenarios compared to prior work [13, 17, 14, 15] (Sections 5.1, 5.2, Appendix A, Appendix B).
In summary, our contributions are:
@@
Our approach is agnostic to the adaptation task (detection or segmentation), target domain (night, bad weather), saliency guidance, model architecture, and domain adaptation algorithm.
@@
Our approach is efficient, with train-time image warping and feature unwarping but no test-time warping, resulting in minimal training overhead and no additional inference latency (Appendix B).
Our approach improves domain adaptation across various lighting (+3.7 mAP50 for BDD100K Day Night detection), weather conditions (+6.1 mAP50 for BDD100K Clear DENSE Foggy detection), and geographies (+6.3 mIoU for Cityscapes ACDC semantic segmentation).
@@ 2 Related Work
Saliency Guided Image Warping. Non-uniform Spatial Transformations have been used to correct image distortions arising from camera acquisition [43]. For example, several spatial transformation methods [16, 24] for learned models introduce intentional image distortions to enhance performance. Learning to Zoom [24] showed that over-sampling task-specific salient regions is useful, as visual recognition requires combining information from different spatial resolutions. Building on this paradigm, warping for other tasks like detection [35, 36] and segmentation [36] has been explored. For instance, Fovea [35] and Two Plane Prior [9] designed ‘salient’ priors to guide warping, introducing static & temporal, and geometric saliency guidance, respectively.
Our approach stands outs in the following ways: (a) Unlike former image warping methods that indiscriminately process object and background, we introduce an analytical instance-level saliency guidance that targets regions containing objects. (b) Unlike previous image warping efforts focused on supervised settings, our novel training strategy and saliency guidance tailor image warping for unsupervised domain adaptation. (c) Unlike prior warping works like [35, 9, 36], our strategy introduces no additional inference latency as we do not warp during testing.
Image Transformations for Domain Adaptation. Image transformations like flipping, cropping, blurring, color jittering, grayscaling, and contrast adjustment are commonly used as perturbations to enhance domain adaptation [38, 46, 47, 23, 29, 49, 37, 6, 41, 44, 18, 13, 14, 12, 15, 17]. These techniques increase source domain diversity and narrow the gap to the target domain. For example, 2PCNet [17] introduces a domain-specific Night Augmentation (NightAug) to make day images resemble night images. To the best of our knowledge, in-place image warping as a domain adaptive perturbation has not been explored previously. Our approach is agnostic to domain adaptation algorithms and can be seamlessly combined with other augmentation strategies.
@@ 3 Method
We first describe how to warp images and unwarp features using a spatial guidance function. Next, we introduce our novel warping guidance to explicitly oversample objects (as shown in Figure 1). Finally, we discuss incorporating our warping strategy into self-training unsupervised domain adaptation algorithms (See Figure 3).
@@ 3.1 Image Warping and Feature Unwarping
We employ the saliency guided sampling mechanism from [24]. However, other methods, such as thin plate spline transformations [16, 7], can also be used.
Image Warping: We consider an inverse transformation parameterized by saliency such that:
@@
Feature Unwarping: If we operate in the space of warped images, our predictions would also be warped. However, we do not wish to warp our labels to match the warped input image, Instead, we perform feature unwarping of spatial backbone features and train the model without warping our labels or unwarping our predictions like [35, 9]. This strategy makes our warping mechanism task-agnostic. We adopt the strategy from LZU [36], which approximates with , a piecewise tiling of invertible bilinear maps, and then perform feature unwarping via:
Additional Considerations. We restrict the space of allowable warps for to preserve axis alignment. In this approach, entire rows or columns are ‘stretched’ or ‘compressed,’ which are shown by [35, 9, 36] to perform better:
where is a Gaussian kernel and represent the two image axes in pixel space. and are the marginalized versions of along and axes, respectively.
@@ 3.2 Designing Saliency Guidance
Saliency () determines which regions of an image are sampled more during warping [24]. We explore three saliency functions: two based on approximate guidance from previous work [35, 36, 9], and a novel instance-level saliency guidance proposed in this work. Since images are warped in-place, expanding some regions necessitates reducing others (Figure 4). Only our instance-level saliency guidance oversamples objects and undersamples the background.
Approximate Saliency Guidance: The Geometric Prior [9] saliency guidance is a function of the dominant vanishing point in the scene, oversampling objects close to the vanishing point, which are usually small due to their distance from the viewpoint. However, regions near the vanishing point may lack object, and not all objects may be sufficiently oversampled (Figure 4).
The Static Prior [35, 36] assumes that object distribution remains consistent at test time, which is often untrue due to domain shift. The incorrect assumption results in a saliency map targeting average object locations, which may fail to accurately represent specific target images where object positions deviate from these averages (Figure 4).
Moreover, warping during test time, after adaptation, not only increases inference latency but is also ineffective (See Section 5.3). Essentially, approximate saliency guidance ignores groundtruth object labels and fails to focus on salient objects, which is crucial for improving adaptation.
Instance-level Saliency Guidance. We propose Instance-level Saliency Guidance which leverages ground truth labels in the source domain (Figure 4). During training, source image has corresponding labels available. For simplicity, we assume the labels are bounding boxes.
Given an image, we aim to oversample object pixels, such as cars, and undersample background pixels, such as roads. Following [1, 35], we use kernel density estimation to model the bounding box saliency as a sum of Gaussians:
@@
If many objects are small, we need to oversample objects more aggressively to mitigate background dominance. Conversely, if objects are already large enough, less aggressive oversampling is required. Therefore, we propose an expansion factor to determine object oversampling intensity:
@@
where is the fraction of object size corresponding to an interval between size thresholds and (See Figure 1). In our case, we consider , to represent the percentage of small, medium, and large objects in the dataset, as per COCO [20] object size thresholds.
@@
Saliency scale and expansion factor are inversely proportional (), where is a constant. We clip between a lower bound of 0, to prevent negative saliency values that lead to holes, and an upper bound , to discourage extreme warping.
3.3 Warping for Unsupervised Domain Adaptation
@@
We consider unsupervised domain adaptation, where the source domain includes labeled data, i.e., , whereas the target domain consists only of unlabeled images, i.e., .
@@
Domain Adaption Framework: The standard framework follows a student-teacher self-training approach. It begins with a supervised pre-training stage where a model is trained using supervised loss on the source domain, Next, in the unsupervised adaptation stage, the model is adapted to the target domain using dataset through an additional unsupervised loss. Generally, a student and teacher (initialized using ) are trained jointly through self-training with psuedo-labels [17, 13]. Following [13, 14, 15, 17], we update the teacher parameters, as DINO [3] and Mean Teachers [34] have shown that this approach leads to better performance compared to traditional distillation. While training the student, a supervised loss is used to maintain performance on the source domain in a supervised branch,
@@
Meanwhile, the teacher is updated as the exponential moving average (EMA) of the student like[13, 14, 15, 17].
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl |
2PCNet | 27.0 | 45.0 | 30.1 | 0.0 | 15.4 | 35.5 |
+ Ours w/ Sta. Prior | 28.6 | 48.5 | 31.7 | 0.0 | 19.6 | 34.0 |
+ Ours w/ Geo. Prior | 27.1 | 45.8 | 29.5 | 0.0 | 15.2 | 37.5 |
+ Ours w/ Inst. Warp | 33.1 | 51.1 | 39.7 | 0.0 | 14.5 | 45.2 |
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl |
2PCNet | 25.8 | 48.4 | 24.4 | 10.9 | 34.1 | 49.0 |
+ Ours w/ Sta. Prior | 26.8 | 50.6 | 24.9 | 11.4 | 35.9 | 49.6 |
+ Ours w/ Geo. Prior | 26.9 | 49.9 | 24.5 | 11.8 | 35.6 | 50.1 |
+ Ours w/ Inst. Warp | 27.1 | 51.4 | 24.6 | 11.0 | 36.7 | 49.0 |
Warping and Unwarping in the Domain Adaptation Framework: Our warping mechanism operates only on source domain images while training, i.e. it is employed in the supervised pre-training stage and in the student model during the unsupervised stage when source images are used. Consider the recognition model as a combination of backbone and prediction heads , i.e. . We warp the image and then unwarp the feature maps from the backbone to compose the final output,
@@
following [36] for a source image . Thus, losses are modified for the supervised pre-training stage as and for the unsupervised training stage as,
We have experimented with warping target images (during training and inference) using Static [35, 36] and Geometric [9] prior saliency guidances, which do not require ground truth labels. However, this approach reduced performance (See Table 8). Finally, we do not warp during test time because it introduces additional latency and our ablation study (Table 8) shows that it does not improve performance after adaptation.
@@
Base Algorithm and Losses: We use DAFormer [13] and 2PCNet [17] as our base segmentation and detection adaption algorithms, respectively. We directly adopt their losses for and without modification. Our method is task-agnostic (segmentation or detection) and makes no additional assumptions about the setup. Therefore, other algorithms [14, 15, 18, 6, 49, 42] are equally applicable.
@@
Inference: We employ the teacher model during inference. Our method does not require image warping or feature unwarping during testing. This contrast with image warping techniques proposed in the supervised setting [35, 9, 36], which incur overheads from test-time warping and require additional information (like vanishing points [9]) to generate saliency maps for warping at test time.
@@
Supervised Setting: Our training method can also be applied in supervised learning scenario by following the supervised pre-training steps to train a model . That is, omitting the unsupervised training stage completely and only performing supervised pre-training stage – then removing the warping and unwarping layers completely from the trained model. This is unlike prior work [24, 35, 9, 36] which also advocates warping at test time. With our warping on the source domain, model performance improves on the target domain even without adaptation (See Table 5 and 6). For Source Domain Results, see Appendix C.
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl |
2PCNet | 23.5 | 46.4 | 21.1 | 9.5 | 26.0 | 43.3 |
+ Ours w/ Sta. Prior | 25.3 | 49.4 | 22.4 | 10.3 | 27.4 | 45.1 |
+ Ours w/ Geo. Prior | 25.6 | 49.7 | 22.9 | 10.0 | 27.7 | 45.7 |
+ Ours w/ Inst. Warp | 25.6 | 50.1 | 22.6 | 9.8 | 27.6 | 46.6 |
Method | mAP | mAP50 | mAP75 | mAPs | mAPm | mAPl |
2PCNet | 16.6 | 30.8 | 14.9 | 6.0 | 24.8 | 26.6 |
+ Ours w/ Sta. Prior | 17.3 | 31.0 | 16.7 | 6.2 | 25.9 | 29.5 |
+ Ours w/ Geo. Prior | 16.6 | 30.3 | 15.4 | 5.2 | 25.1 | 28.7 |
+ Ours w/ Inst. Warp | 17.9 | 32.1 | 18.2 | 6.8 | 26.7 | 32.1 |